1. A man has a rise in income from 25000 to 27000; as a result his savings rise from 900 to 1300. What is his MPC?
  2. As income increases from 15000 to 17000 and savings from 5000 to 5800. What is MPC?
  3. If in an economy, $20 is taxed out of every additional $100 in national income, $20 is saved and $10 is spent on imports. What is the value of the multiplier?
  4. If in an economy, out of every additional 2000 of NY , 400 is taxed, 450 is saved and 100 is spent on imports. What is the value of the multiplier?
  5. An economy has a NY OF 65m, a MPS of 0.3 and a MPT of 0.1. If the government wishes to achieve a NNY of 75m, how much will they have to increase spending by?
  6. A man has an increase income from 35000 to 38000, as result his consumption increased from 14000 to 20000. What is his MPS?
1
Expert's answer
2021-11-11T17:22:35-0500

1) The marginal propensity to consume is equal to ΔC / ΔY, where ΔC is the change in consumption, and ΔY is the change in income. If consumption increases by 80 cents for each additional dollar of income, then MPC is equal to 0.8 / 1 = 0.8.

ΔY=2700-2500=2000

consumption increases 2000-(1300-900)=2000-400=1600

ΔC=1600

"MPC=\frac{1600}{2000}=\frac{4}{5}=\frac{0.8}{1}=0.8"

2) ΔY=17000-15000=2000

consumption increases 2000 - (5800-5000)=2000-800=1200

ΔC=1600

"MPC=\frac{1200}{2000}=\frac{3}{5}=\frac{0.6}{1}=0.6"